# NaturalSpeech 3: Zero-Shot Speech Synthesis with Factorized Codec and Diffusion Models

### v1

@@ Abstract
While recent large-scale text-to-speech (TTS) models have achieved significant progress, they still fall short in speech quality, similarity, and prosody. Considering speech intricately encompasses various attributes (e.g., content, prosody, timbre, and acoustic details) that pose significant challenges for generation, a natural idea is to factorize speech into individual subspaces representing different attributes and generate them individually. Motivated by it, we propose NaturalSpeech 3, a TTS system with novel factorized diffusion models to generate natural speech in a zero-shot way. Specifically, 1) we design a neural codec with factorized vector quantization (FVQ) to disentangle speech waveform into subspaces of content, prosody, timbre, and acoustic details; 2) we propose a factorized diffusion model to generate attributes in each subspace following its corresponding prompt. With this factorization design, NaturalSpeech 3 can effectively and efficiently model the intricate speech with disentangled subspaces in a divide-and-conquer way. Experiments show that NaturalSpeech 3 outperforms the state-of-the-art TTS systems on quality, similarity, prosody, and intelligibility. Furthermore, we achieve better performance by scaling to 1B parameters and 200K hours of training data.
@@ 1 Introduction
Benefiting from these designs, NaturalSpeech 3 has achieved significant improvements in speech quality, similarity, prosody, and intelligibility. Specifically, 1) it achieves comparable or better speech quality than the ground-truth speech on the LibriSpeech test set in terms of CMOS; 2) it achieves a new SOTA on the similarity between the synthesized speech and the prompt speech (0.64 $\rightarrow$ 0.67 on Sim-O, 3.69 $\rightarrow$ 4.01 on SMOS); 3) it shows a significant improvement in prosody compared to other TTS systems with --0.16 average MCD (lower is better), $+$0.21 SMOS; 4) it achieves a SOTA on intelligibility (1.94 $\rightarrow$ 1.81 on WER). Furthermore, we demonstrate the scalability of NaturalSpeech 3 by scaling it to 1B parameters and 200K hours of training data. Audio samples can be found in https://speechresearch.github.io/naturalspeech3.
@@ 3.2.1 FACodec Model Overview
We propose a factorized neural speech codec (i.e., FACodec111 We will release the code and pre-trained checkpoint of FACodec soon.) to convert complex speech waveform into disentangled subspaces representing speech attributes of content, prosody, timbre, and acoustic details and reconstruct high-quality speech waveform from these.
@@ 3.2.2 Attribute Disentanglement
Figure 3: The framework of factorized diffusion model, which consists of 1) phoneme encoder, 2) duration diffusion and length regulator, 3) prosody diffusion, 4) content diffusion, 5) detail (acoustic detail) diffusion. Note that modules 2-5 shares the same diffusion formulation.
@@ 3.3.2 Diffusion Formulation
Inference. During inference, we progressively replace masked tokens, starting from the fully masked sequence $\mathbf{X}_{T}$, by iteratively sampling from $p(\mathbf{X}_{t-\Delta t}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$. Inspire by [50, 51, 52], we first sample $\mathbf{\hat{X}}_{0}$ from $p_{\theta}(\mathbf{X}_{0}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$, and then sample $\mathbf{X}_{t-\Delta t}$ from $q(\mathbf{X}_{t-\Delta t}|\mathbf{\hat{X}}_{0},\mathbf{X}_{t})$, which involves remask $\lfloor N\cdot\sigma(t-\Delta t)\rfloor$ tokens in $\mathbf{\hat{X}}_{0}$ with the lowest confidence score, where we define the confidence score of $\hat{x}_{i}$ in $\mathbf{\hat{X}}_{0}$ to $p_{\theta}(\hat{x}_{i}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$ if $m_{t,i}=0$, otherwise, we set confidence score of $x_{i}$ to $1$, which means that tokens already unmasked in $\mathcal{X}_{t}$ will not be remasked.
@@ 3.4 Connections to the NaturalSpeech Series
  • [leftmargin=*]
  • Goal. The NaturalSpeech series aims to generate natural speech with high quality and diversity. We approach this goal in several stages: 1) Achieving high-quality speech synthesis in single-speaker scenarios. To this end, NaturalSpeech [4] generates speech with quality on par with human recordings and only tackles single-speaker recording-studio datasets (e.g., LJSpeech). 2) Achieving high-quality and diverse speech synthesis on multi-style, multi-speaker, and multi-lingual scenarios. Both NaturalSpeech 2 [5] and NaturalSpeech 3 focus on speech diversity by exploring the zero-shot synthesis ability based on large-scale, multi-speaker, and in-the-wild datasets.
@@ 4.1 Experimental Settings
Table 1: The evaluation results for NaturalSpeech 3 and the baseline methods on LibriSpeech test-clean. ${}^{\spadesuit}$ means the results are obtained from the authors. means the results directly obtained from the paper. ${}^{\clubsuit}$ means the results are infered from offical checkpoints. means the reproduced results. Abbreviation: LT (LibriTTS), V (VCTK), LJ (LJSpeech), ${}^{\star}$ (Librilight Small, Medium), EX (Expresso), MS (MSSS Kor), NI (NIKL Kor).
@@
NaturalSpeech ${}^{\spadesuit}$ Librilight 0.55 0.62 1.94 -0.18 3.65
${}^{\spadesuit}$ Self-Collected (60kh) 0.64 0.67 2.03 -0.23 3.69
@@
Mega-TTS ${}^{\spadesuit}$ Librilight 0.53 - 2.32 -0.20 3.63
${}^{\spadesuit}$ Mixed (165kh) 0.57 0.68 2.49 -0.25 3.71
StyleTTS ${}^{\clubsuit}$ LT + V + LJ 0.38 - 2.49 -0.21 3.07
${}^{\clubsuit}$ LT + ${}^{\star}$ + EX + MS + NI 0.51 - 6.33 -0.41 3.50
@@
Evaluation Metrics. Objective Metrics: In the Librispeech test-clean benchmark, we evaluate both speaker-similarity (SIM-O and SIM-R) and robustness (WER). In specific, 1) for SIM-O and SIM-R, we employ the WavLM-TDCNN [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification speaker embedding model to assess speaker similarity between generated samples and the prompt. Results are reported for both similarity to original prompt (SIM-O) and reconstructed prompt (SIM-R). 2) for Word Error Rate (WER), we use an ASR model [^3] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe generated speech. The model is a CTC-based HuBERT pre-trained on Librilight and fine-tuned on the 960 hours training set of LibriSpeech. In the RAVDESS benchmark, we evaluate the prosody similariy (MCD and MCD-Acc). In specific, 1) following [59], we adopt Mel-Ceptral Distortion (MCD) for prosody evaluation by measuring the differences between generated samples and ground truth samples. We report the results for eight emotions, along with the average result. 2) for MCD-Acc, we evaluate the top-1 emotion accuracy of the generated speech on the RAVDESS benchmark for prosodic similarity measures. Specifically, we adopt a K-Nearest-Neighbors (KNN) model as emotion classifier. We compare MCD distances between the generated speech and the ground-truth speech from the same speaker, across eight different emotions. Subjective Metrics: We employ comparative mean option score (CMOS) and similarity mean option score (SMOS) in both two benchmarks to evaluate naturalness and similarity, respectively.
@@ 4.2.2 Generation Similarity
Table 2: The evaluation results for NaturalSpeech 3 and the baseline methods on RAVDESS. ${}^{\spadesuit}$ means the results are obtained from the authors. ${}^{\clubsuit}$ means the results are inferred from official checkpoints. means the reproduced results. Abbreviation: Avg (average MCD), Acc (MCD-Acc).
@@
Mega-TTS ${}^{\spadesuit}$ 4.44 0.39 -0.20 4.51
StyleTTS ${}^{\clubsuit}$ 4.50 0.40 -0.25 3.98
${}^{\clubsuit}$ 6.08 0.30 -0.37 3.87
@@ 4.3.1 Ablation Study
Table 5: The reconstruction quality evaluation of codecs. ${}^{\clubsuit}$ means results are infered from offical checkpoints. ${}^{\bigstar}$ means the reproduced checkpoint. means the reproduced model following the original paper’s implementation and experimental setup. All models use a codebook size of $1024$. Bold for the best result and underline for the second-best result. Abbreviation: H (Hop Size), N (Codebook Number).
@@
${}^{\clubsuit}$ 24kHz 320 8 6.0 kbps 3.28 0.94 0.99 2.70
${}^{\clubsuit}$ 16kHz 320 4 2.0 kbps 3.17 0.93 0.98 3.05
${}^{\clubsuit}$ 16kHz 320 9 4.5 kbps 3.52 0.95 0.97 2.65
@@ 4.3.2 Method Analyses
Table 6: The comparison between autoregressive approach with (VALL-E + F) and without (VALL-E) our proposed factorization on LibriSpeech test-clean. means the reproduced results. Abbreviation: Sim-O/R (Sim-O / Sim-R).
@@
Extensibility. NaturalSpeech 3 utilizes a non-autoregressive model for discrete token generation with factorization design. To validate the extensibility of our proposed factorization method, we further explore the autoregressive generative model for discrete token generation under our factorization framework. We utilize VALL-E for verification. We first employ an autoregressive language model to generate prosody codes, followed by a non-autoregressive model to generate the remaining content and acoustic details codes. This approach maintains a consistent order of attribute generation, allowing for a fair comparison. We name it VALL-E + F. As shown in Table 6, VALL-E + F consistently outperforms VALL-E by a considerable margin in all objective and subjective metrics, demonstrating the factorization design can enhance VALL-E in speech similarity, quality and generation robustness. It further shows our factorization paradigm is not limited in the proposed factorization diffusion model and has a large potential in other generative models. We leave it for future work.
@@ 5 Conclusion
In this paper, we develop a TTS system that consists of 1) a novel neural speech codec with factorized vector quantization (i.e., FACodec) to decompose speech waveform into distinct subspaces of content, prosody, acoustic details and timbre and 2) novel factorized diffusion model to synthesize speech by generating attributes in subspaces with discrete diffusion. NaturalSpeech 3 outperforms the state-of-art TTS system on speech quality, similarity, prosody, and intelligibility. We also show that NaturalSpeech 3 can enable speech attribute manipulation, by customizing speech attribute prompts. Furthermore, we demonstrate that NaturalSpeech 3 achieves better performance by scaling to 1B parameters and 200K hours of training data. We list the limitations and future works in Appendix C.
@@ A.3 Evaluation Baselines
A.4 Latency Analysis
@@
A.5 Ablation Study on Duration Diffusion Model
@@
A.6 Details of Prosody Similarity Evaluation
@@
Table 11: The MCD scores on $8$ different emotions of NaturalSpeech 3 and the baseline methods on RAVDESS. ${}^{\spadesuit}$ means the results are obtained from the authors. ${}^{\clubsuit}$ means the results are inferred from official checkpoints. means the reproduced results. We use bold to indicate the best result and underline to indicate the second-best result.
@@
NaturalSpeech ${}^{\spadesuit}$ 2.77 3.51 4.85 4.88 5.42 5.23 5.31 4.52
Mega-TTS ${}^{\spadesuit}$ 3.28 4.39 4.44 4.67 4.21 5.00 5.42 4.14
StyleTTS ${}^{\clubsuit}$ 3.41 4.38 4.40 4.64 4.80 4.69 5.10 4.57
${}^{\clubsuit}$ 5.54 6.55 5.78 5.84 6.37 6.17 6.74 5.62

### v2

@@ Abstract
While recent large-scale text-to-speech (TTS) models have achieved significant progress, they still fall short in speech quality, similarity, and prosody. Considering speech intricately encompasses various attributes (e.g., content, prosody, timbre, and acoustic details) that pose significant challenges for generation, a natural idea is to factorize speech into individual subspaces representing different attributes and generate them individually. Motivated by it, we propose NaturalSpeech 3, a TTS system with novel factorized diffusion models to generate natural speech in a zero-shot way. Specifically, 1) we design a neural codec with factorized vector quantization (FVQ) to disentangle speech waveform into subspaces of content, prosody, timbre, and acoustic details; 2) we propose a factorized diffusion model to generate attributes in each subspace following its corresponding prompt. With this factorization design, NaturalSpeech 3 can effectively and efficiently model the intricate speech with disentangled subspaces in a divide-and-conquer way. Experiments show that NaturalSpeech 3 outperforms the state-of-the-art TTS systems on quality, similarity, prosody, and intelligibility, and achieves on-par quality with human recordings. Furthermore, we achieve better performance by scaling to 1B parameters and 200K hours of training data.
@@ 1 Introduction
Benefiting from these designs, NaturalSpeech 3 has achieved significant improvements in speech quality, similarity, prosody, and intelligibility. Specifically, 1) it achieves comparable or better speech quality than the ground-truth speech on the LibriSpeech test set in terms of CMOS; 2) it achieves a new SOTA on the similarity between the synthesized speech and the prompt speech (0.64 $\rightarrow$ 0.67 on Sim-O, 3.69 $\rightarrow$ 4.01 on SMOS); 3) it shows a significant improvement in prosody compared to other TTS systems with --0.16 average MCD (lower is better), $+$0.21 SMOS; 4) it achieves a SOTA on intelligibility (1.94 $\rightarrow$ 1.81 on WER); 5) it achieves human-level naturalness on multi-speaker datasets (e.g., LibriSpeech), another breakthrough after NaturalSpeech [^1] While NaturalSpeech 1 [4] achieved human-level quality on the single-speaker LJSpeech dataset, NaturalSpeech 3 achieved human-level quality on the diverse multi-speaker LibriSpeech dataset for the first time.. Furthermore, we demonstrate the scalability of NaturalSpeech 3 by scaling it to 1B parameters and 200K hours of training data. Audio samples can be found in https://speechresearch.github.io/naturalspeech3.
@@ 3.2.1 FACodec Model Overview
We propose a factorized neural speech codec (i.e., FACodec222 We will release the code and pre-trained checkpoint of FACodec soon.) to convert complex speech waveform into disentangled subspaces representing speech attributes of content, prosody, timbre, and acoustic details and reconstruct high-quality speech waveform from these.
@@ 3.2.2 Attribute Disentanglement
Figure 3: The framework of factorized diffusion model, which consists of 1) phoneme encoder, 2) duration diffusion and length regulator, 3) prosody diffusion, 4) content diffusion, 5) detail (acoustic detail) diffusion. Note that modules 2-5 share the same diffusion formulation.
@@ 3.3.2 Diffusion Formulation
Inference. During inference, we progressively replace masked tokens, starting from the fully masked sequence $\mathbf{X}_{T}$, by iteratively sampling from $p(\mathbf{X}_{t-\Delta t}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$. Inspire by [50, 51, 52], we first sample $\mathbf{\hat{X}}_{0}$ from $p_{\theta}(\mathbf{X}_{0}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$, and then sample $\mathbf{X}_{t-\Delta t}$ from $q(\mathbf{X}_{t-\Delta t}|\mathbf{\hat{X}}_{0},\mathbf{X}_{t})$, which involves remask $\lfloor N\cdot\sigma(t-\Delta t)\rfloor$ tokens in $\mathbf{\hat{X}}_{0}$ with the lowest confidence score, where we define the confidence score of $\hat{x}_{i}$ in $\mathbf{\hat{X}}_{0}$ to $p_{\theta}(\hat{x}_{i}|\mathbf{X}_{t},\mathbf{X}^{p},\mathbf{C})$ if $m_{t,i}=1$, otherwise, we set confidence score of $x_{i}$ to $1$, which means that tokens already unmasked in $\mathcal{X}_{t}$ will not be remasked.
@@ 3.4 Connections to the NaturalSpeech Series
  • Goal. The NaturalSpeech series aims to generate natural speech with high quality and diversity. We approach this goal in several stages: 1) Achieving high-quality speech synthesis in single-speaker scenarios. To this end, NaturalSpeech [4] generates speech with quality on par with human recordings and only tackles single-speaker recording-studio datasets (e.g., LJSpeech). 2) Achieving high-quality and diverse speech synthesis on multi-style, multi-speaker, and multi-lingual scenarios. NaturalSpeech 2 [5] firstly focuses on speech diversity by exploring the zero-shot synthesis ability based on large-scale, multi-speaker, and in-the-wild datasets. Furthermore, NaturalSpeech 3 further achieves human-level naturalness on the multi-speaker dataset (e.g., LibriSpeech).
@@ 4.1 Experimental Settings
Evaluation Metrics. Objective Metrics: In the Librispeech test-clean benchmark, we evaluate speaker-similarity (SIM-O and SIM-R), speech quality (UTMOS), and robustness (WER). In specific, 1) for SIM-O and SIM-R, we employ the WavLM-TDCNN [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification speaker embedding model to assess speaker similarity between generated samples and the prompt. Results are reported for both similarity to original prompt (SIM-O) and reconstructed prompt (SIM-R); 2) for speech quality, we employ UTMOS [59] which is a surrogate objective metric of MOS; 3) for Word Error Rate (WER), we use an ASR model [^4] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe generated speech. The model is a CTC-based HuBERT pre-trained on Librilight and fine-tuned on the 960 hours training set of LibriSpeech. We also use an advanced ASR model based on transducer [60] [^5] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge. In the RAVDESS benchmark, we evaluate the prosody similarity (MCD and MCD-Acc). In specific, 1) following [61], we adopt Mel-Ceptral Distortion (MCD) for prosody evaluation by measuring the differences between generated samples and ground truth samples. We report the results for eight emotions, along with the average result. 2) for MCD-Acc, we evaluate the top-1 emotion accuracy of the generated speech on the RAVDESS benchmark for prosodic similarity measures. Specifically, we adopt a K-Nearest-Neighbors (KNN) model as emotion classifier. We compare MCD distances between the generated speech and the ground-truth speech from the same speaker, across eight different emotions. Subjective Metrics: We employ comparative mean option score (CMOS) and similarity mean option score (SMOS) in both two benchmarks to evaluate naturalness and similarity, respectively.
Table 1: The evaluation results for NaturalSpeech 3 and the baseline methods on LibriSpeech test-clean. means the results are obtained from the authors. means the results directly obtained from the paper. means the results are infered from offical checkpoints. means the reproduced results. Abbreviation: LT (LibriTTS), V (VCTK), LJ (LJSpeech), LL (Librilight Small, Medium), EX (Expresso), MS (MSSS Kor), NI (NIKL Kor). Please refer to Appendix A.4 for more results on 1) WER inferred by an advanced ASR system, and 2) UTMOS, an automatic metric for MOS.
@@
NaturalSpeech 2 Librilight 0.55 0.62 1.94 -0.18 3.65
Voicebox Self-Collected (60kh) 0.64 0.67 2.03 -0.23 3.69
@@
Mega-TTS 2 Librilight 0.53 - 2.32 -0.20 3.63
UniAudio Mixed (165kh) 0.57 0.68 2.49 -0.25 3.71
StyleTTS 2 LT + V + LJ 0.38 - 2.49 -0.21 3.07
HierSpeech++ LT + LL + EX + MS + NI 0.51 - 6.33 -0.41 3.50
@@ 4.2.2 Generation Similarity
Table 2: The evaluation results for NaturalSpeech 3 and the baseline methods on RAVDESS. means the results are obtained from the authors. means the results are inferred from official checkpoints. means the reproduced results. Abbreviation: Avg (average MCD), Acc (MCD-Acc).
@@
Mega-TTS 2 4.44 0.39 -0.20 4.51
StyleTTS 2 4.50 0.40 -0.25 3.98
HierSpeech++ 6.08 0.30 -0.37 3.87
@@ 4.2.3 Robustness
4.2.4 Human-Level Naturalness on LibriSpeech Testset
We compare the speech synthesized by NaturalSpeech 3 with human recordings (Ground Truth) in Table 1 (more results can be found in Table 9 in Appendix A.4). We have the following observations: 1) NaturalSpeech 3 achieves -0.01 Sim-O and +0.16 SMOS compared to human recordings, which demonstrates that our method is on par or better on speaker similarity; 2) NaturalSpeech 3 achieves -0.08 CMOS and +0.16 UTMOS compared with recording, which demonstrates that our method can generate on-par or better voice quality; 3) Our method also achieves close WER with human recordings, which demonstrates the robustness of NaturalSpeech 3. Therefore, we can conclude that for the first time, NaturalSpeech 3 has achieved human-level quality and naturalness on the multi-speaker LibriSpeech test set in a zero-shot way. It is another great milestone after NaturalSpeech 1 [4] has achieved human-level quality on the single-speaker LJSpeech dataset.
@@ 4.3.1 Ablation Study
Table 5: The reconstruction quality evaluation of codecs. means results are infered from offical checkpoints. means the reproduced checkpoint. means the reproduced model following the original paper’s implementation and experimental setup. All models use a codebook size of $1024$. Bold for the best result and underline for the second-best result. Abbreviation: H (Hop Size), N (Codebook Number).
@@
EnCodec 24kHz 320 8 6.0 kbps 3.28 0.94 0.99 2.70
HiFi-Codec 16kHz 320 4 2.0 kbps 3.17 0.93 0.98 3.05
DAC 16kHz 320 9 4.5 kbps 3.52 0.95 0.97 2.65
@@ 4.3.2 Method Analyses
Table 6: The comparison between autoregressive approach with (VALL-E + FACodec) and without (VALL-E) our proposed factorization on LibriSpeech test-clean. means the reproduced results. Abbreviation: Sim-O/R (Sim-O / Sim-R).
@@
Extensibility. NaturalSpeech 3 utilizes a non-autoregressive model for discrete token generation with factorization design. To validate the extensibility of our proposed factorization method, we further explore the autoregressive generative model for discrete token generation under our factorization framework. We utilize VALL-E for verification. We first employ an autoregressive language model to generate prosody codes, followed by a non-autoregressive model to generate the remaining content and acoustic details codes. This approach maintains a consistent order of attribute generation, allowing for a fair comparison. We name it VALL-E + FACodec. As shown in Table 6, VALL-E + FACodec consistently outperforms VALL-E by a considerable margin in all objective and subjective metrics, demonstrating the factorization design can enhance VALL-E in speech similarity, quality and generation robustness. It further shows our factorization paradigm is not limited in the proposed factorization diffusion model and has a large potential in other generative models. We leave it for future work.
@@ 5 Conclusion
In this paper, we develop a TTS system that consists of 1) a novel neural speech codec with factorized vector quantization (i.e., FACodec) to decompose speech waveform into distinct subspaces of content, prosody, acoustic details and timbre and 2) novel factorized diffusion model to synthesize speech by generating attributes in subspaces with discrete diffusion. NaturalSpeech 3 outperforms the state-of-the-art TTS system on speech quality, similarity, prosody, and intelligibility. We also show that NaturalSpeech 3 can enable speech attribute manipulation, by customizing speech attribute prompts. Furthermore, we demonstrate that NaturalSpeech 3 achieves human-level performance on the multi-speaker LibriSpeech dataset for the first time and better performance by scaling to 1B parameters and 200K hours of training data. We list the limitations and future works in Appendix C.
@@ A.3 Evaluation Baselines
A.4 More Experimental Results on Zero-shot TTS
In this section, we report more evaluation results for NaturalSpeech 3 and other baselines on: 1) WER, inferred by an advanced ASR system [^9] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge; 2) UTMOS [59], which is a surrogate objective metric of MOS. The results are shown in Table 9.
Table 9: The evaluation results for NaturalSpeech 3 and the baseline methods on LibriSpeech test-clean. means the results are obtained from the authors. means the results directly obtained from the paper. means the results are inferred from offical checkpoints. means the reproduced results. WER means the word error rate calculated by an advanced ASR system mentioned in  A.4.
[…]
Ground Truth 0.68 - 1.94 0.68 4.14 +0.08 3.85
VALL-E - 0.58 5.90 - - - -
VALL-E 0.47 0.51 6.11 4.87 3.68 -0.60 3.46
NaturalSpeech 2 0.55 0.62 1.94 1.24 3.88 -0.18 3.65
Voicebox 0.64 0.67 2.03 1.81 3.82 -0.23 3.69
Voicebox 0.48 0.50 2.14 1.24 3.73 -0.32 3.52
Mega-TTS 2 0.53 - 2.32 2.17 4.02 -0.20 3.63
UniAudio 0.57 0.68 2.49 1.81 3.79 -0.25 3.71
StyleTTS 2 0.38 - 2.49 1.58 3.94 -0.21 3.07
HierSpeech++ 0.51 - 6.33 4.97 3.80 -0.41 3.50
NaturalSpeech 3 0.67 0.76 1.81 1.13 4.30  0.00 4.01
A.5 Latency Analysis
@@
A.6 Ablation Study on Duration Diffusion Model
@@
A.7 Details of Prosody Similarity Evaluation
@@
Table 12: The MCD scores on $8$ different emotions of NaturalSpeech 3 and the baseline methods on RAVDESS. means the results are obtained from the authors. means the results are inferred from official checkpoints. means the reproduced results. We use bold to indicate the best result and underline to indicate the second-best result.
@@
NaturalSpeech 2 2.77 3.51 4.85 4.88 5.42 5.23 5.31 4.52
Mega-TTS 2 3.28 4.39 4.44 4.67 4.21 5.00 5.42 4.14
StyleTTS 2 3.41 4.38 4.40 4.64 4.80 4.69 5.10 4.57
HierSpeech++ 5.54 6.55 5.78 5.84 6.37 6.17 6.74 5.62